kofi is four times as old as kobby and kofi is 36years older than kobby. how old is kofi?
Answer by ikleyn(52943)

.
Answer.   Kobby is  12  years old;  Kofi is  48  years old.

Actually, this problem can be solved mentally and momentarily, as soon as you completed reading the condition.


But if you want to use Algebra, here it is:


    4x - x = 36

    3x = 36

    x = 36/3 = 12  is the Kobby age.